OpenStudy (anonymous):

lim 2x³-x²+7x-3 = x->∞ 2-x+5x²-4x³

OpenStudy (solomonzelman):

http://www.wolframalpha.com/

ganeshie8 (ganeshie8):

\[\large \lim \limits_{x\to \infty}~ \dfrac{2x^3-x^2+7x-3}{2-x+5x^2-4x^3}\]

ganeshie8 (ganeshie8):

like this ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

divide top and bottom by x^3 as \(x\to \infty, \frac{1}{x}\to 0 \)

ganeshie8 (ganeshie8):

\[\lim \limits_{x\to \infty}~ \dfrac{2x^3-x^2+7x-3}{2-x+5x^2-4x^3} = \lim \limits_{\frac{1}{x}\to 0}~ \dfrac{2-\frac{1}{x}+\frac{7}{x^2}-\frac{3}{x^3}}{\frac{2}{x^3}-\frac{1}{x^2}+\frac{5}{x}-4} \\ = \dfrac{2-0+0-0}{0-0+0-4} = \dfrac{-1}{2}\]

OpenStudy (anonymous):

I caught \[2x^{3} and -4x^{3} = -1/2\]

ganeshie8 (ganeshie8):

taking the ratio of leading coefficients works too^

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

yw

OpenStudy (anonymous):

can help me more one question

OpenStudy (anonymous):

\[\frac{ x }{ \sqrt{x+1} }\]

OpenStudy (anonymous):

limite infinity positive